# Yintong Huo‡§, Michael R. Lyu

### v1

@@ VI-E2 Component-based Implementation Limitation
Figure 8: A webpage with multiple elements of the same structure and style.
@@
10 <img src="./placeholder.svg" alt="ManUtdvsAthleticBilbao" class="w-full" />
@@
Finding 8: MLLMs show critical deficiencies in component-based implementation and framework-specific syntax, revealing fundamental limitations in producing reusable front-end code.
VI-E3 Issue Detection Limitation
In Design Repair task, we also prompt MLLMs to judge the UI display issues described in Section IV-C. Then we calculate the issue detection accuracy by comparing the MLLMs’ outputs with the ground truth annotated by human experts. Table IX presents the UI issue detection accuracy across different MLLM models. The results demonstrate consistently poor performance of MLLMs in identifying UI issues, with average 0.2972, 0.2205 and 0.2275, 0.3403 rate on React, Vue, Angular and Vanilla, respectively.
TABLE IX: UI Issue Identification Accuracy of different MLLMs.
@@ VI-E3 Issue Detection Limitation
VII Threats to Validity
@@
Internal validity. (1) The scores generated by MLLM-as-a-judge may have reliability concerns. To eliminate this, we carefully design the prompt and provide the model with a detailed guideline of evaluation criteria. The results of human evaluation in Section V-B also verified that the accuracy of MLLM’s score is above 90%. (2) Potential data leakage. Half of the webpages in our benchmark are sourced from closed applications (e.g., top 500 websites by Moz [37]), and the ground truth for design repair tasks is manually written by developers, minimizing data leakage risks. Moreover, the poor framework-based code generation performance, evident in syntax errors and limited use of framework-specific features, suggests that MLLM results in the benchmark are unlikely due to mere data memorization.
@@
External validity. We only include limited frameworks of React, Vue, and Angular, due to their dominance in modern web development. These frameworks collectively represent the majority of contemporary applications and offer diverse programming paradigms: from React’s JSX to Vue’s templates and Angular’s TypeScript approach, providing comprehensive evaluation coverage.
VIII Discussion
We elicit several actionable advice from our findings for researchers and developers in this field.
@@ VIII Discussion
For researchers: (1) Enhance framework-specific MLLM training. The poor performance on framework-specific syntax and component-based implementations underscores the need for MLLM training datasets enriched with modern web development patterns and framework-specific best practices. This would improve their practical value for applying to diverse front-end ecosystems. (2) Improve multimodal information fusion for UI tasks. Our findings reveal that for design edit and repair tasks, single code input and multimodal input perform equally well, suggesting that MLLMs currently underutilize visual information. Future work should prioritize developing more effective visual-code alignment and specialized attention mechanisms for multimodal reasoning in front-end jobs.
@@
For developers: (1) Provide code edit location information for Design Edit and Design Repair tasks. Since MLLMs struggle with code localization, providing exact locations for edits or repairs would significantly enhance performance and reduce the cognitive burden on models to identify relevant code segments. (2) Clearly state the repair issues for repair tasks. Given MLLMs’ low accuracy in UI issue identification (27%), explicitly stating the problems allows models to focus on targeted solutions rather than inaccurate independent diagnosis. (3) Decompose complex instructions and large designs. Given that MLLM performance degrades with complex instructions and large UI mockups, practitioners can improve the practicality by breaking down complex requirements into simpler, atomic tasks and segmenting large UI designs into smaller, manageable components. This approach transforms challenging tasks into more tractable problems that align better with current MLLM capabilities.
IX Conclusion

### v2

@@ 6.5.2. Component-based Implementation Limitation
Figure 8. A webpage with multiple elements of the same structure and style.
@@
10 <img src="./place.svg" alt="ManUtdvsAthleticBilbao" class="w-full" />
@@
Finding 8: MLLMs show critical deficiencies in component-based implementation and framework-specific syntax, revealing limitations in producing reusable front-end code.
6.5.3. Issue Detection Limitation
In Design Repair task, we also prompt MLLMs to judge the UI display issues described in Section 4.3. Then we calculate the issue detection accuracy by comparing the MLLMs’ outputs with the ground truth annotated by human experts. Table 10 presents the UI issue detection accuracy across different MLLM models. The results demonstrate consistently poor performance of MLLMs in identifying UI issues, with average 0.2972, 0.2205 and 0.2275, 0.3403 rate on React, Vue, Angular and Vanilla, respectively.
Table 10. UI Issue Identification Accuracy of different MLLMs.
@@ 6.5.3. Issue Detection Limitation
6.6. RQ6: What mistakes do MLLMs make on the three tasks?
We conduct systematic error analysis by randomly selecting 25% of results for examination. The annotation process involves iterative discussion and refinement of failure categories until consensus is reached, with continuous communication to identify emerging failure types during annotation.
For the design generation task, we define seven failure categories: (1) Compile error: generated code cannot be compiled successfully; (2) Layout disorder: UI arrangement deviates from the reference design; (3) Element missing: essential UI components are absent; (4-7) Element attribute errors: wrong position, color, text content, or size of UI elements. For design edit and repair tasks, we establish six failure types: (1) Compile error: modified code contains compilation errors; (2) No edit/repair: models make no modifications; (3) Wrong objects: models target incorrect elements; (4) Wrong edit/repair: models apply inappropriate modifications to correct elements; (5) Partial edit/repair: incomplete implementation of required changes; (6) Unnecessary modification: extraneous adjustments beyond specified requirements. The detailed annotation guideline and failure type examples are in our repo [^4] https://github.com/WebPAI/DesignBench/blob/main/assets/Failure_Annotation_Guideline.md.
(a) Design Generation.
(b) Design Edit.
(c) Design Repair.
Figure 9. Failure type distribution on three tasks of different MLLMs.
Finding 10: MLLMs exhibit distinct task-specific failure patterns reflecting different cognitive demands. Design generation fails primarily on spatial reasoning (wrong element size/position) and structural completeness (element missing), design editing suffers from scope control issues (unnecessary modifications, partial edits), while design repair shows fundamental limitations in issue element identification and solution formulation (high rates of no repair attempts).
From Figure 9, we observe distinct failure patterns across tasks: (1) Design generation tasks are dominated by spatial and structural failures, with ”wrong element size,” ”element missing,” and ”wrong element position” as the top three failure types, revealing MLLMs’ limitations in layout comprehension and spatial reasoning. (2) Design edit tasks show more controlled failure modes, with ”unnecessary modifications,” ”partial edits,” and ”wrong edits” as primary concerns, indicating challenges in precise instruction following and modification scope control. (3) Design repair tasks exhibit the most severe limitations, with ”no repair,” ”wrong repairs,” and ”partial repairs” as dominant failure types, demonstrating fundamental difficulties in defect identification and corrective action formulation.
7. Threats to Validity
@@
Internal validity. (1) The MLLM-as-a-judge scores may have reliability concerns. To eliminate this, we carefully design the prompt and provide the model with a detailed guideline of evaluation criteria. The results of human evaluation in Section 5.4 also verified that the accuracy of MLLM’s score is above 90%. (2) Potential data leakage. Half of the webpages in our benchmark are sourced from closed applications (e.g., top 500 websites by Moz (41)), and the ground truth for design repair tasks is manually written by developers, minimizing data leakage risks. Moreover, the poor framework-based code generation performance, evident in syntax errors and limited use of framework-specific features, suggests that MLLM results in the benchmark are unlikely due to mere data memorization. We also demonstrate that there is a low risk of data contamination in Section 5.3.
@@
External validity. We only include limited frameworks of React, Vue, and Angular, due to their dominance in modern web development. These frameworks represent the majority of contemporary applications and offer diverse programming paradigms: from React’s JSX to Vue’s templates and Angular’s TypeScript approach, providing comprehensive evaluation coverage.
8. Discussion
@@
We elicit several actionable advice from our findings for researchers and developers in this field. (1) For researchers: (i) Enhance framework-specific MLLM training. The poor performance on framework-specific syntax and component-based implementations underscores the need for MLLM training datasets enriched with modern web development patterns and framework-specific best practices. This would improve their practical value for applying to diverse front-end ecosystems. (ii) Improve multimodal information fusion for UI tasks. Our findings reveal that for design edit and repair tasks, single code input and multimodal input perform equally well, suggesting that MLLMs currently underutilize visual information. Future work should prioritize developing more effective visual-code alignment for multimodal reasoning in front-end jobs.
@@
(2) For developers: (i) Provide code edit location information for Design Edit and Design Repair tasks. Since MLLMs struggle with code localization, providing exact locations for edits or repairs would significantly enhance performance and reduce the cognitive burden on models to identify relevant code segments. (ii) Clearly state the repair issues for repair tasks. Given MLLMs’ low accuracy in UI issue identification, explicitly stating the problems allows models to focus on targeted solutions rather than inaccurate independent diagnosis. (iii) Decompose complex instructions and large designs. Given that MLLM performance degrades with complex instructions and large UI mockups, practitioners can improve the practicality by breaking down complex requirements into simpler, atomic tasks and segmenting large UI designs into smaller, manageable components.
9. Conclusion